Circular No. 7380

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000Q IN ANONYMOUS GALAXY Y. Sato and W. D. Li, University of California at Berkeley (UCB), on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of a supernova (mag ∼ 17.8) on an unfiltered image taken on Mar. 12.5 UT. SN 2000Q is located at α = 16^{h}05^{m}11^{s}.85, δ = +69^{o}39′51′′.5 (equinox 2000.0), which is 4′′.1 east and 1′′.6 north of the nucleus of an anonymous galaxy that is in turn ∼ 218′′ east and 7′′ south of the galaxy NGC 6079. An unfiltered image taken on Mar. 1.5 showed nothing at the position of the new object (limiting mag ∼ 19.0). A. V. Filippenko and A. L. Coil, UCB, report that uncalibrated CCD spectra (range 330–1000 nm), obtained on Mar. 15 with the Shane 3-m reflector at Lick Observatory, reveal that the supernova is of type Ia, possibly subluminous, with quite prominent Si II features at rest wavelengths of 590 and 615 nm. The very approximate recession velocity derived from the Si II observed wavelengths is 6000 km/s.

SUPERNOVA 2000P IN NGC 4965 E. Cappellaro, S. Benetti, M. Turatto, and A. Pastorello, Padua, re-

port that two spectra (range 384–684 nm, resolution 0.7 nm) of SN 2000P (cf. IAUC 7378, 7379), obtained on Mar. 11.2 and 14.3 UT with the Danish 1.54-m telescope (+ DFOSC) at the European Southern Observatory, confirm that this object is a type-II supernova close to maximum. Indeed, the spectrum shows H Balmer lines overimposed on a blue continuum. The profiles of Balmer lines (Hα, Hβ, Hγ, Hδ, Hϵ) show at least two P-Cyg components: a narrow one with absorption minima corresponding to an expansion velocity of ∼ 500 km/s and a broader component with expansion velocity ∼ 8000 km/s (FWHM), in analogy to SNe 1996L and 1994aj (Benetti et al. 1999, MNRAS 305, 811; 1998, ibid. 294, 448). Narrow P-Cyg components of similar width are visible also for Fe II (multiplet 42) and Ca II H and K.

NO NEW TRANSIENT AX J1747.1–2845 H. Murakami et al. retract their item on IAUC 7379, saying that AX

J1747.1–2845 appears to be merely a false image of SAX J1747.0-2853 itself, caused by improper processing with the quick-look software.

2000 March 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green